# Negative impact of heavy-tailed uncertainty and error distributions on the reliability of calibration statistics for machine learning regression tasks

### v3

@@ V.2 Comparison of validation scores
The statistics and bootstrapped confidence intervals have been estimated for the RCE and ZMS for all datasets with $10^{4}$ bootstrap replicates. The reference values $\vartheta_{ref}$ are 0 for the RCE and 1 for the ZMS. The resulting values and validation scores are reported in Table 4. It is clear from these results that average calibration is not satisfied by all datasets and, more problematically, that the diagnostic depends strongly on the choice of statistic.
@@
Globally, the statistics disagree on more than half of the datasets, which is surprising for two statistics deriving analytically from the same generative model (Eq. 1). Considering that the RCE does not involve the pairing of uncertainties and errors implicit in the ZMS, one could have expected it to be more forbidding, which is the case for Sets 2, 3 and 4, but not for Set 6 (maybe also Sets 5 and 7).
It is remarkable that the five sets for which a disagreement between RCE and ZMS is observed (Sets 2, 3, 4, 6 and 7) are those with the largest skewness values for $u_{E}^{2}$ (Table 3, Fig. 5), with $\beta_{GM}(u_{E}^{2})\geq 0.6$, except for Set 6, as shown in Fig. 5(left). This suggests that the right tail of the uncertainty distribution plays a major role in this disagreement. Sets 2, 4, 6 and 7 present also $\beta_{GM}(E^{2})\geq 0.8$, which might point to problematic error distributions.
The role of the tails of the uncertainty and error distributions is assessed in the next section.
V.3 Role of the distributions tails
In order to better understand the discrepancy of the validation results by RCE and ZMS, one needs to consider the sensitivity of these statistics to the uncertainty distributions, and notably to the large, sometimes outlying, values, as suggested by the skewness and kurtosis analyses. In the z-scores, these large uncertainty values are likely to contribute to small absolute values of $Z$ having a small impact on the ZMS, while they are likely to affect significantly the estimation of the RCE. This hypothesis is tested on the nine datasets by a decimation experiment, where both statistics are estimated on datasets iteratively pruned from their largest uncertainties.
@@ V.3 Role of the distributions tails
It appears that in all cases ZMS is less or as sensitive as RCE to the large extremal uncertainty values and that its decimation curve always strays within the limits of the corresponding CI. For RCE, one can find cases where it is more sensitive than ZMS but lies within the limits of the CI (Sets 2 and 6) and cases where it strays beyond the limits of the CI (Sets 3, 4 and 7). These five cases are precisely those where the RCE diagnostic differs the most from the ZMS (Sect. V.2). This sensitivity test confirms the hypothesis that RCE is more sensitive than ZMS to the upper tail of the uncertainty distribution, to a point where its estimation might become unreliable.
For some sets (2, 3, 6 and 7) a small percentage (1-4 %) of the largest uncertainties affects the $\Delta_{RCE}$ statistic, after which one observes a sharp change of slope, and the curves becomes a reflected image of the $\Delta_{ZMS}$ statistic. These points might therefore be considered as outlying values. By contrast, one observes for Set 4 a smooth transition, which suggest that the shape of the tail is involved, more than a few extreme uncertainty values.
@@
Surprisingly, the $\zeta_{RCE}$ values for Sets 2 and 6 do not follow the same trend, with a deterioration for Set 2 and no notable change for Set 6. The main reason for the discrepancy of validation metrics for Sets 2 and 6 is therefore not the shape of the uncertainty distributions, which leads us to the error distributions. Note that these are the sets for which the $\Delta_{RCE}$ curve does not stray notably out of the corresponding CI limits. Both datasets present very large error skewness values $\beta_{GM}(E^{2})$ (0.92 and 0.96) and kurtosis values $\kappa_{CS}$ (19.7 and 22.7), well above their respective safety limits. The skewness of z-scores $\beta_{GM}(Z^{2})$ for these sets (0.83 and 0.95), although smaller than $\beta_{GM}(E^{2})$, are also above the safety limit. In comparison, the kurtosis for $\kappa_{CS}(Z^{2})$ for Set 2 (6.4) is much smaller than for Set 6 (24.0), while still above the safety limit (5.0). It is therefore difficult to conclude on the average calibration of Sets 2 and 6, as both the RCE and ZMS are likely to be affected by the heavy-tails of the $E^{2}$ and $Z^{2}$ distributions, respectively.
V.4 Impact of binning on tailedness metrics
The previous findings for average calibration statistics might also impact the reliability of conditional calibration. The UCE and ENCE statistics are also based on the comparison of MV to MSE , but, being bin-based, they might be less susceptible to be affected by this sensitivity issue when the binning variable is $u_{E}$. For a large enough number of bins, the problem of heavy tails or large outlying $u_{E}$ values should vanish for individual bins. However, this is not necessarily the case for the corresponding error and z-scores subsets, and one might still be left with unreliable binned RCE statistics.
To assess the impact of $u_{E}$-binning on the tailedness statistics $\beta_{GM}$ and $\kappa_{CS}$ of the binned data, they were estimated for a set of 20 equal-sized bins and reported in Fig.8. The bins number is chosen to avoid small bins which might cause additional statistical reliability problems. The boxplots show the dispersion of the statistics of the 20 bins of $u_{E}^{2}$, $E^{2}$ and $Z^{2}$, while the red points locate the corresponding statistic for the full dataset.
Figure 8: Impact of binning on the tailedness parameters of the binned data (20 bins). The red points locate the corresponding statistic for the full dataset. The red dashed line is the safety limit provided in Table 2.
As expected for the binned $u_{E}^{2}$ variable, the improvement of both $\beta_{GM}$ and $\kappa_{CS}$ is remarkable, and all statistics get below the safety limit (red dashed line), except for a few rare outliers. Skewness lies around zero, and kurtosis gets negative values, reflecting the nearly rectangular or truncated shape of the binned $u_{E}^{2}$ distributions. A notable exception is Set 7, which has stratified uncertainty values resulting from an isotonic regression calibration treatment(Pernot2023b_arXiv, ). The splitting of such a comb-like distribution into equal-size bins might result into pathological discrete distributions, as evidenced from the boxplots. This binning problem has been explored in a previous study(Pernot2023d, ).
Concerning the binned errors, one can see that the effect on $\beta_{GM}(E^{2})$ is set-dependent and that there is no overall trend. Some binned datasets get an improvement and get below the safety limit (Sets 3 and 4), others get an improvement but stay at least partly above the safety limit (Sets 2, 5 and 7) and some do not get any notable improvement (Sets 5, 6, 8 and 9), or even a degradation as some bins get over the safety limit (Sets 5 and 8). The observations are globally similar for kurtosis, with again a very large dispersion for Set 7, aggravated by the occurrence of an infinite kurtosis value (not shown).
The z-scores do not benefit from a noticeable improvement for any set, and for all datasets the tailedness values get dispersed around the full set value. In consequence, even for a priori non-problematic datasets, a non-negligible part of the bins might rise above the safety limits. The slight improvement that can be seen on skewness for Sets 1, 3, 4 and 5 is not observed for kurtosis.
To summarize, unless the tailedness of $u_{E}^{2}$ is the sole problem for a dataset, one cannot expect $u_{E}$ binning to alleviate the reliability problem for binned calibration statistics. Moreover, ff the UCE, ENCE or LZMS analysis are based on another binning variable than uncertainty (such as an input feature), as is requested to test adaptivity(Pernot2023d, ), there is no reason for an improvement of the binned distributions for any of $u_{E}^{2}$, $E^{2}$ and $Z^{2}$. In consequence, if a dataset has excessive tailedness statistics, there is no fundamental reason for conditional calibration to be more reliable than average calibration.
@@ VI Conclusions
Discrepancies of validation diagnostic was observed between both statistics for an ensemble of datasets extracted from the recent ML-UQ literature for regression tasks. This anomaly has been elucidated by showing that the RCE is very sensitive to the upper tail of the uncertainty distribution, and notably to the presence of outliers, which has been proven by a decimation experiment (5 cases out of 9). Moreover, error sets with heavy tails can also be challenging for the RCE (5 cases). The distribution of z-scores ($Z=E/u_{E}$) appeared to be less problematic, although some heavy-tailed $Z$ distributions have been observed (2 cases). The underlying problem is our inability to design reliable confidence intervals with a prescribed coverage for the mean of variables with heavy-tailed distributions ($u_{E}^{2}$, $E^{2}$ and $Z^{2}$). Bootstrapping is presently the best approach to CI estimation for the mean of non-normal variables, and numerical experiments have shown that it fails in conditions representative of the studied datasets. A major consequence of these observations is that average calibration statistics based on the comparison of $MSE=<E^{2}>$ to $MV=<u_{E}^{2}>$ should not be blindly relied upon for the kind of datasets found in ML-UQ regression problems. The RCE is affected by both MSE and MV sensitivity to heavy tails. In contrast, the ZMS statistic has fewer reliability issues and should therefore be the statistic of choice for average calibration testing.
@@
In this context, it is important to be able to screen out problematic datasets. It has been shown that robust skewness ($\beta_{GM}$) and/or kurtosis ($\kappa_{CS}$) statistics can be used for this, and safety thresholds have been defined for both statistics and for the $u_{E}^{2}$, $E^{2}$ and $Z^{2}$ distributions. The reliability of the RCE and ZMS statistics for datasets with values exceeding these limits has to be questioned.
@@
A temptation would be to ignore altogether average calibration statistics and to focus on conditional calibration. As the RCE, the popular UCE and ENCE statistics implement the comparison of MV to MSE, but, being bin-based, one might expect them to be less susceptible to the above-mentioned sensitivity issue when the binning variable is $u_{E}$ (consistency testing(Pernot2023d, )). Unfortunately, it has be shown in this study that even if binning solves the reliability problem for $<u_{E}^{2}>$, this is not the case for $<E^{2}>$, leaving these conditional calibration statistics with the same reliability problem as for the RCE. This also applies to local ZMS analysis. It is therefore essential to evaluate the tailedness of the distributions before any calibration assessment.
@@
In post hoc calibration, scaling factors for the uncertainties can be derived from the ZMS ($\sigma$ scaling(Laves2020, ), BVS(Frenkel2023, ; Pernot2023c_arXiv, )) or from the UCE(Frenkel2023, ). So, by extension, this problem also affects post-hoc calibration procedures based on statistics related to the RCE, ENCE or ZMS (for instance the NLL), and those parameterized on datasets with heavy uncertainty and/or error tails might also be unreliable.
@@
In the absence of ML-UQ methods designed to mitigate this heavy tails problem [^1] Note that uncertainty sets with large dispersion (large $c_{v}$ statistic) and therefore potentially heavy tails, are favored for active learning., there is no solution to safely ensure and validate the calibration of uncertainties for heavy-tailed uncertainty and error distributions. For such datasets, the best available approach is to rely on graphical methods(Pernot2022a, ; Pernot2022b, ) and to derive a qualitative diagnostic. A change of uncertainty paradigm, using prediction intervals or distributions instead of standard deviations, seems presently the most promising alternative to recover a sound statistical validation framework for calibration.
@@ Acknowledgments
The code and data to reproduce the results of this article are available at https://github.com/ppernot/2024_RCE/releases/tag/v2.0 and at Zenodo (https://doi.org/10.5281/zenodo.10998413).

### v4

@@ V.2 Comparison of validation scores
The statistics and bootstrapped confidence intervals have been estimated for the RCE and ZMS for all datasets with $10^{4}$ bootstrap replicates. The resulting values and validation scores are reported in Table 4. It is clear from these results that average calibration is not satisfied by all datasets and, more problematically, that the diagnostic depends notably on the choice of statistic.
@@
Globally, the statistics disagree on more than half of the datasets, which is surprising for two statistics deriving analytically from the same generative model (Eq. 1). Considering that the RCE ignores the pairing of uncertainties and errors, one could have expected it to be more forbidding than the ZMS, which is the case for Sets 2, 3 and 4, but not for Set 6 (maybe also for Sets 5 and 7).
It is remarkable that the five sets for which a disagreement between RCE and ZMS is observed (Sets 2, 3, 4, 6 and 7) are those with the largest skewness values for $u_{E}^{2}$ (Table 3, Fig. 6), with $\beta_{GM}(u_{E}^{2})\geq 0.6$, except for Set 6, as shown in Fig. 6(left). This suggests that the upper tail of the uncertainty distribution plays a major role in this disagreement. Sets 2, 4, 6 and 7 present also $\beta_{GM}(E^{2})\geq 0.8$, which might point to problematic error distributions.
The impact of the tails of the uncertainty and error distributions is assessed in the next section.
V.3 Impact of the distributions tails
In order to better understand the discrepancy of the validation results by RCE and ZMS, one needs to consider the sensitivity of these statistics to the uncertainty distributions, and notably to the large, sometimes outlying, values, as suggested by the skewness analysis. In the z-scores, these large uncertainty values are likely to contribute to small absolute values of $Z$ having a small impact on the ZMS, while they are likely to affect significantly the estimation of the RCE. This hypothesis is tested on the nine datasets by a decimation experiment, where both statistics are estimated on datasets iteratively pruned from their largest uncertainties.
@@ V.3 Impact of the distributions tails
It appears that in all the cases ZMS is less or as sensitive as RCE to the large extremal uncertainty values and that its decimation curve always strays within the limits of the corresponding CI. For RCE, one can find cases where it is more sensitive than ZMS but lies within the limits of the CI (Sets 2 and 6) and cases where it strays beyond the limits of the CI (Sets 3, 4 and 7). These five cases are precisely those where the RCE diagnostic differs the most from the ZMS (Sect. V.2). This sensitivity test confirms the hypothesis that RCE is more sensitive than ZMS to the upper tail of the uncertainty distribution, to a point where its estimation might become unreliable.
For some sets (2, 3, 6 and 7) a small percentage (1-4 %) of the largest uncertainties affects the $\Delta_{RCE}$ statistic, after which one observes a sharp change of slope, and the curves becomes a reflected image of the $\Delta_{ZMS}$ statistic. These uncertainties might therefore be considered as outlying values. By contrast, one observes for Set 4 a smooth transition, which suggest that the shape of the tail is involved, rather than a few extreme uncertainty values.
@@
Surprisingly, the $\zeta_{RCE}$ values for Sets 2 and 6 do not follow the same trend, with a deterioration for Set 2 and no notable change for Set 6. Note that these are the sets for which the $\Delta_{RCE}$ curve does not stray notably out of the corresponding CI limits. The main reason for the discrepancy of validation metrics for Sets 2 and 6 is therefore not the shape of the uncertainty distributions, which leads us to the error distributions.
Both datasets present very large error skewness values $\beta_{GM}(E^{2})$ (0.92 and 0.96) , well above the safety limit of 0.8. The skewness of z-scores $\beta_{GM}(Z^{2})$ for these sets (0.83 and 0.95), although smaller than $\beta_{GM}(E^{2})$, are also above the safety limit. It is therefore difficult to conclude on the average calibration of Sets 2 and 6, as both the RCE and ZMS are likely to be affected by the heavy-tails of the $E^{2}$ and $Z^{2}$ distributions, respectively.
@@ VI Conclusions
Discrepancies of validation diagnostic was observed between both statistics for an ensemble of datasets extracted from the recent ML-UQ literature for regression tasks in materials and chemical sciences. This anomaly has been elucidated by showing that the RCE is very sensitive to the upper tail of the uncertainty distribution, and notably to the presence of outliers, which has been proven by a decimation experiment (5 cases out of 9). Moreover, error sets with heavy tails can also be challenging for the RCE (5 cases). The distribution of z-scores ($Z=E/u_{E}$) appeared to be less problematic, although some heavy-tailed $Z$ distributions have been observed (2 cases).
Two underlying problems contribute to this state of affairs: (1) strong sensitivity of the Mean Squares statistic to the presence of heavy-tails and outliers; and (2) our inability to design reliable confidence intervals with a prescribed coverage for the mean of variables with heavy-tailed distributions ($u_{E}^{2}$, $E^{2}$ and $Z^{2}$). Bootstrapping is presently the best approach to CI estimation for the mean of non-normal variables, and numerical experiments have shown that it fails in conditions representative of the studied datasets.
A major consequence of these observations is that average calibration statistics based on the comparison of $MSE=<E^{2}>$ to $MV=<u_{E}^{2}>$ should not be blindly relied upon for the kind of datasets found in ML-UQ regression problems. The RCE is affected by both MSE and MV sensitivity to heavy tails. In contrast, the ZMS statistic has fewer reliability issues and should therefore be the statistic of choice for average calibration testing.
@@
In this context, it is important to be able to screen out problematic datasets. It has been shown that a robust skewness statistic ($\beta_{GM}$) can be used for this, and safety thresholds have been defined for the $u_{E}^{2}$, $E^{2}$ and $Z^{2}$ distributions. The reliability of the RCE and ZMS statistics for datasets with values exceeding these limits has to be questioned.
@@
A temptation would be to ignore altogether average calibration statistics and to focus on conditional calibration. As the RCE, the popular UCE and ENCE statistics implement the comparison of MV to MSE, but, being bin-based, one might expect them to be less susceptible to the above-mentioned sensitivity issue when the binning variable is $u_{E}$ (consistency testing(Pernot2023d, )). Unfortunately, even if binning solves the tailedness problem for $<u_{E}^{2}>$ for most of the bins, this is not the case for $<E^{2}>$, leaving these conditional calibration statistics with the same reliability problem as the RCE. This also applies to the local ZMS analysis. The problem of conditional calibration statistics will be further explored in a forthcoming article.
@@
In post hoc calibration, scaling factors for the uncertainties can be derived from the ZMS ($\sigma$ scaling(Laves2020, ), BVS(Frenkel2023, ; Pernot2023c_arXiv, )) or from the UCE(Frenkel2023, ). So, by extension, the outlined problems also affect post-hoc calibration procedures based on statistics related to the RCE, ENCE or ZMS (for instance the NLL). Those methods parameterized on datasets with heavy uncertainty and/or error tails might also be unreliable.
@@
One can envision three paths to get out of this lack of reliability: (1) the reduction of the problematic tails of uncertainty and error distributions can be handled to some extent by iterative learning, i.e. by iteratively feeding the data with the largest absolute errors and uncertainties to the training algorithm until the problem subsides(Pernot2020b, ); (2) for datasets with uncontrollable heavy tails, one might have to abandon formal validation and to rely on graphical methods to derive a qualitative calibration diagnostic(Pernot2022a, ; Pernot2022b, ); and (3) a change of uncertainty paradigm, using prediction intervals or distributions instead of standard deviations, seems presently to be the most promising alternative to recover a sound statistical validation framework for calibration.
@@ Acknowledgments
The code and data to reproduce the results of this article are available at https://github.com/ppernot/2024_RCE/releases/tag/v4.0 and at Zenodo (https://doi.org/10.5281/zenodo.11489178).
Appendix A An alternative formulation of RCE
The biased estimation of the RCE as defined by Eq. II.1 for heavy-tailed uncertainty and error distributions can be reduced by expressing it as
$RCE2=\frac{MV-MSE}{MV}$ (22)
where the mean variance (MV) and mean squared errors (MSE) are used instead of their square roots.
Simulations with the same setup as in Sect. IV.2 were performed for the RCE2 statistic. The results are reported in Fig. 9.
Figure 9: Estimated values of an alternative RCE formulation for a series of datasets generated by TIG models and characterized by the error skewness parameter $\beta_{GM}(E^{2})$.
When compared to the results for the original RCE [Fig. 4(left)], two features are remarkable: (1) the mean values present also deviations increasing with $\beta_{GM}$, but with positive or negative signs; and (2) the Monte Carlo $2\sigma$ error bars now cover the reference value, so that there is no systematic bias for RCE2.